OpenStudy (anonymous):

Find the slope and y-intercept of the line. y=1/5x-8? please help!!!!

TheSmartOne (thesmartone):

You are missing something. :)

OpenStudy (anonymous):

any one ?? <:3

TheSmartOne (thesmartone):

Like an x.

TheSmartOne (thesmartone):

Is it y=1/5x-8 ??

TheSmartOne (thesmartone):

@derpstar We need to know your question because there is a typo...

TheSmartOne (thesmartone):

@derpstar Are you there?

OpenStudy (anonymous):

its 1/5x not 1/5 whoops

TheSmartOne (thesmartone):

Ok. Well you see the question is in the format of y=mx+b where m is the slpe and b is the y-intercept. So y=mx+b y=1/5x-8 What does m=? and b=? m will give you the answer for the slope and b will tell you the y-intercept.

OpenStudy (anonymous):

oooooh this is super easy m=1/5 b=8 correct??

TheSmartOne (thesmartone):

m=1/5 is correct

TheSmartOne (thesmartone):

@dmndlife24 Do not give answers.

OpenStudy (anonymous):

yeah thanks guys and i forgot the negative lol but thanks so much for the help really appreciate it!!!!

TheSmartOne (thesmartone):

You see it is y=mx \(\Huge + \) b y=1/5x \(\Huge -\) 8

TheSmartOne (thesmartone):

No prob

TheSmartOne (thesmartone):

lol I helped her and you just gave her the answer... *shakes head* oh well